# Channel Measurement, Modeling, and Simulation for 6G: A Survey and Tutorial

### v3

@@ III THz Channel Measurements and Models
As presented in Section I, 6G uses a higher frequency spectrum than previous generations [6] to achieve multi-Gbps data rate services such as immersive communication. Recently, new frequency bands beyond 5G, such as mid bands [71] and THz bands [72], have been researched to meet the requirement of the large data rates. Among these frequency bands, THz has been identified as a promising spectrum in 6G due to the wide bandwidth and low interference characteristics, and the ITU has shown strong support for the research and development of THz technologies. As shown in Fig. 4, the ITU has allocated a total of 234.2 GHz in the 100-450 GHz frequency band for fixed services and land mobile applications. Specifically, 12 bands (total 97.2 GHz bandwidth) within the 100-275 GHz frequency range [40] have been allocated by ITU. Of note, the largest contiguous bandwidth available in the frequency bands is 23 GHz. Below 200 GHz, the largest contiguous bandwidth currently allocated for fixed or mobile applications is 12.5 GHz. The limited availability of contiguous bandwidths poses a challenge for developing high-bandwidth wireless communication systems in 6G. Therefore, by the ITU Radio Regulations 2020, the WRC 2019 (WRC-19) identified 275-296 GHz, 306-313 GHz, 318-333 GHz, and 356-450 GHz, for a total of 137 GHz, which can be used for fixed service and land mobile applications [73]. Moreover, the maximum available continuous bandwidth is 94 GHz.
@@
The WRC is not the only organization planning for THz spectrum. Also, spectrum planning and allocation regulators in different countries and regions have released parts of the THz band to facilitate scientific research. For example, Japan’s Ministry of Information and Communication (MIC) designated the frequency range of 116-134 GHz as “Commercial Telecommunications Services” in 2015, and the 152-164 GHz and 287.5-312.5 GHz bands were made available for experimental testing in 2020 [74]. In 2018, the European Conference of Postal and Telecommunications Administrations (CEPT) and the European Radiocommunications Committee (ERC) released a recommendation on “short range devices,” also known as “unlicensed devices,” which specified provisions for these devices to operate within the frequency ranges of 122-122.25 GHz and 244-246 GHz [75]. The Federal Communications Commission (FCC) decision in March 2019 authorized the unlicensed use of 116-123 GHz, 174.8-182 GHz, 185-190 GHz, and 244-246 GHz in the United States [76]. In conclusion, a high priority has been placed on THz spectrum development and utilization in various countries and organizations.
@@
III-A THz Channel Measurements
@@
It is important to conduct extensive THz measurements to build channel model for THz communication [77]. Up to now, many research organizations and institutions conduct multiple THz channel measurement campaigns in the abovementioned frequency bands. We have classified and summarized some representative THz channel measurement campaigns by frequency bands, striving to comprehensively cover all currently measured THz frequency bands (100-1100 GHz), as shown in Table I.
@@
In the THz spectrum allocated for land mobile applications, as illustrated in Fig. 4, THz measurement campaigns are relatively concentrated around the frequencies of 100, 140, 220, and 300 GHz [78, 17, 26, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122, 123, 124, 125, 98, 126, 127, 128, 129, 130, 131, 132, 133, 134, 109]. The frequency of 100 GHz represents the upper limit of the ITU channel model. The primary objective of most studies is to investigate the applicability of existing channel models in the THz frequency band. Due to the non-linear factors of atmospheric absorption, the selection of actual frequencies primarily considers the “atmospheric window” frequency bands. In the THz frequency band, the “atmospheric window” refers to special frequency bands, such as 140 GHz and 220 GHz, where propagation attenuation is relatively low. Hence, these bands have been the focus of extensive research. Additionally, 300 GHz falls within the lower boundary of the THz frequency range defined by the ITU radio regulations, making it a frequency band of significant interest. However, in the 450-1100 GHz band [130, 135, 136, 137], there are fewer channel measurement campaigns limited by capabilities of the hardware, such as instruments and equipment. Due to limitations in transmitted power, the majority of studies focus only on reflection, scattering, and attenuation characteristics.
@@
Secondly, the measurement platforms in the summarized THz channel measurement campaigns are introduced. Most measurement platforms widely used in THz channel measurement campaigns are frequency domain vector network analyzer (VNA)-based platform [138]. In the VNA-based platform, the Tx and Rx are directly connected to the VNA. The simple structure, the synchronized Tx and Rx, the large bandwidth, and the high dynamic range make this measurement platform occupy the mainstream. However, due to the high loss of the cable, the measurement distance is usually short but can be extended to 100 m by radio-over-fiber (RoF) [109]. To conduct long-distance channel measurements, correlation-based time-domain channel measurement platform is used, which operates with the strong autocorrelation of pseudo-random noise (PN) sequence. It has also been used in several THz channel measurements. Since Tx and Rx are independent of each other and do not require cable connections, the platform is preferable for measurements in outdoor environments [79]. In addition, THz time-domain spectroscopy system (THz-TDS)-based [139] channel measurement platforms are limited by the output power and are commonly used for measurements of reflection and diffraction properties of materials.
@@
Thirdly, the scenarios of the channel campaigns are presented. THz band channel measurements focus on desktops, meeting rooms, offices, data centers, computer motherboards, and indoor factory scenarios in the indoor scenarios, and UMi, canyon, and street in the outdoor scenarios, covering vision scenarios, such as wireless local area networks (WLAN), wireless personal area networks (WPAN), data center networks (DCN), and intradevice communication (IDC) [140]. However, the wireless backhaul/fronthaul scenario has been poorly studied [141]. Most of the channel measurement campaigns in these scenarios are in the 100-450 GHz frequency bands due to the limited hardware performance in the high-frequency bands. Furthermore, studies include the channel characteristics, which will be described in detail in the next subsection.
TABLE I: Summary of main THz channel measurement campaigns
@@
Indoor
PL, SF, DS
@@
corridor)
PL, SF, DS, AS, K, cluster numbers, intra/inter-cluster parameters [117, 118]
@@
III-B Important Channel Propagation Characteristics in THz Bands
Current THz characteristics research can be mainly categorized into two types [26]. The first focuses on the material characteristics at THz bands, which actually have an impact on the radio propagation (e.g., reflection) [142]. The second is dedicated to newly emerging channel characteristics in THz bands, such as channel sparsity.
III-B1 Reflection characteristics
Research on propagation mechanism characteristics is the key foundation and fundamental premise for THz channel characteristics research. The reflection are fundamental propagation mechanism that should be thoroughly characterized to establish possible THz wireless links due to the short wavelength introduced in Section I. Therefore, this subsection focuses on the analysis and modeling of the reflection propagation characteristics of THz waves.
@@ III-B1 Reflection characteristics
At low frequencies, the surface of the material is considered primarily as a smooth surface. However, as the frequency increases to the THz band, the wavelength is short enough to be very close to the surface roughness of typical materials. Therefore, in the THz band, the material cannot be considered a smooth surface. The modified reflection coefficient is obtained by introducing a roughness factor $\rho$. The Rayleigh roughness factor can be expressed as [145]
@@
where $\sigma_{h}$ denotes the standard deviation height of the material surface roughness. Then, the modified reflection coefficient $\Gamma^{\prime}$ is expressed as
@@
The modified reflection coefficient model is a function of frequency and angle of incidence, and in reference [79], this phenomenon was verified by experiments. As an example, based on the modified Fresnel model, the frequency-angle two-dimensional reflection coefficient modeling for vertical polarization is proposed. The model of non-metallic materials is written as:
@@
where $a$, $b$, $c$, $d$ are the parameters to be fitted, and their means are shown in [79]. The results for a plasterboard are shown in Fig. 5. The analysis revealed a low root-mean-square (RMS) error which is around 0.1 for the materials of glass, tile, board, and plasterboard, indicating a fit between the points of measured data and the modified reflection coefficient model. As observed, the reflection coefficient of the material and its growth rate increased with increasing angle of incidence. Therefore, the model characterizes the reflection coefficient as a function of the frequency and angle of incidence.
@@
III-B2 Path loss
THz signals transmitted over channels have distinct characteristics compared to lower frequencies, leading to novel and urgent challenges [43]. Representatively, in wireless propagation, the spreading loss increases quadratically with frequency. Besides, with the THz band operating at exceptionally high frequencies, signals transmitted over a THz link are severely affected by the spreading loss [17]. The channel characteristics of high propagation loss can be characterized by path loss. A novel index is essential for characterizing and modeling the path loss in THz channel. Modeling of path loss in the THz band follows the 5G modeling methods, such as the FI model, ABG model, and CI model shown in (1), (3), and (2), respectively, in Section II.A. Based on the references in Table I, the PLE of the CI model $\alpha_{\mathrm{CI}}$ for indoor and outdoor scenarios is summarized, as shown in Fig. 6. The PLE of the THz band is close to 2 in most cases, especially in LOS scenarios. In addition, the PLEs of the NLOS scenarios are markedly higher than those of the LOS scenario, even up to 4.6. This is because the blocking losses caused by walls, cars, etc., can be as large as tens of dB in NLOS scenarios, which means that severe attenuation can occur.
@@ III-B2 Path loss
Due to the wide span of the THz band, it will show a stronger frequency dependence. To further investigate the frequency dependence of path loss in THz bands, [17] used the ABG model to fit the measured path loss. The measured path loss and ABG model are plotted in Fig. 7. The frequency dependence factors $\gamma_{\mathrm{ABG}}$ and distance dependence factors $\alpha_{\mathrm{ABG}}$ are 2.1 and 1.93, respectively. This denotes that the frequency dependence is slightly stronger than that in the free space [43].
@@
Figure 7: Multiple-frequency measurement data and path loss modeling results from 220 to 330 GHz [17].
III-B3 Channel sparsity

### v4

@@ V Multi-Frequency Channel Measurements and Models
As presented in Section I, 6G uses a higher frequency spectrum than previous generations [6] to achieve multi-Gbps data rate services such as immersive communication. Recently, new frequency bands beyond 5G, such as mid-band [136] and THz-band [137], have been researched to meet the requirements of the high data rates. Among these frequency bands, the new WRC cycle will focus on the following bands for the IMT: 4400-4800 MHz, 7125-8400 MHz, and 14.8-15.35 GHz. In December 2023, the 3GPP TSG RAN Rel-19 discussed the research work in 7–24 GHz band. In addition, the THz band has been identified as a promising spectrum in 6G due to the wide bandwidth and low interference characteristics, and the ITU has shown strong support for the research and development of THz technologies. As shown in Fig. 14, the ITU has allocated a total of 234.2 GHz in the 100-450 GHz frequency band for fixed services and land mobile applications. Specifically, 12 bands (total 97.2 GHz bandwidth) within the 100-275 GHz frequency range [53] have been allocated by the ITU. Notably, the largest contiguous bandwidth available in the frequency bands is 23 GHz. Below 200 GHz, the largest contiguous bandwidth currently allocated for fixed or mobile applications is 12.5 GHz. The limited availability of contiguous bandwidths poses a challenge for developing high-bandwidth wireless communication systems in 6G. Therefore, according to the ITU Radio Regulations 2020, the WRC 2019 (WRC-19) identified 275-296 GHz, 306-313 GHz, 318-333 GHz, and 356-450 GHz, for a total of 137 GHz, which can be used for fixed service and land mobile applications [138]. Moreover, the maximum available continuous bandwidth is 94 GHz.
@@
The WRC is not the only organization planning for the THz spectrum. Additionally, spectrum planning and allocation regulators in different countries and regions have released parts of the THz band to facilitate scientific research. For example, Japan’s Ministry of Information and Communication (MIC) designated the frequency range of 116-134 GHz as “Commercial Telecommunications Services” in 2015, and the 152-164 GHz and 287.5-312.5 GHz bands were made available for experimental testing in 2020 [139]. In 2018, the European Conference of Postal and Telecommunications Administrations (CEPT) and the European Radiocommunications Committee (ERC) released a recommendation on “short range devices,” also known as “unlicensed devices,” which specified provisions for these devices to operate within the frequency ranges of 122-122.25 GHz and 244-246 GHz [140]. The Federal Communications Commission (FCC) decision in March 2019 authorized the unlicensed use of 116-123 GHz, 174.8-182 GHz, 185-190 GHz, and 244-246 GHz in the United States [141]. In conclusion, high priority has been placed on THz spectrum development and utilization in various countries and organizations.
@@
V-A New Mid-band Channel Characteristics
In [142], a comprehensive summary of multi-scenario and multi-frequency channel characteristics based on new mid-band (6-24 GHz) channel measurements, including urban macrocell (UMa), urban microcell (UMi), and O2I, was presented. In [136], the presence of some obstacles (e.g., walls, trees, street light poles, and vehicles) in the environment caused the PLEs in the LoS condition to exceed the FSPL model ($n$ = 2). Nevertheless, these values still align well with the PLE ($n$ = 2.1) of UMi-LoS in the 3GPP. A comparison of the PLEs across various LoS conditions and frequency bands reveals that the PLEs are greater in the NLoS condition than in the LoS condition and that they further increase with increasing frequency. This observation indicates that path loss increases at a faster rate with increasing frequency in the NLoS condition.
Under LoS conditions, although the PLE difference among the four bands is very small, the shadow fading of the 15 GHz band is relatively low. Under NLoS conditions, the shadow fading in the 15 GHz band is relatively low, indicating that the rate of increase in path loss is more comparable than that in the 3.3 GHz band, thus demonstrating good coverage capability. Furthermore, it is obvious that although the coverage capability at 6.5 GHz and 15 GHz is clearly weaker than that at 3.3 GHz, these frequency bands are more capable of achieving stronger signal coverage than the 28 GHz band.
The extensive UMi outdoor propagation measurement campaign at 6.75 and 16.95 GHz was conducted via a 1 GHz bandwidth sliding correlation channel sounder[143]. The path loss index, RMS DS and the mean value of RMS AS were obtained. The authors of [65] conducted channel propagation measurements ranging from 2 to 73.5 GHz in three typical outdoor scenarios: UMi Street Canyon, UMi open-air Square, and UMa. The results show that, compared with the three-parameter ABG model, the goodness of fit of the single-parameter CI model in the LoS and NLoS environments is very similar.
Figure 15: The photography of UMi scenario. (a) Layout of UMi scenario. (b) Transceiver antennas location.[136]
Figure 16: Frequency dependence model for mean of delay spread.[144]
Fig. 15 and Fig. 16 show the distributions of the RMS delay spread probability density in the 3.3, 6.5, 15 and 28 GHz frequency bands in the UMi scenario. Fig 16 shows that the higher the frequency is, the smaller the delay. This is because as the frequency increases, the path loss also increases, resulting in a significant attenuation of multipath power. Furthermore, as the carrier frequency increases, the multipath density decreases and the degree of dispersion increases. Therefore, the channel is controlled by a few effective paths, and most of the energy of the MPCs is concentrated in a relatively small area of the delay domain. In the UMi scenario, the RMS delay spread of 28 GHz is approximately several nanoseconds, which is consistent with the measurement results given in [145]. The average RMS delay spread obtained in the low-frequency band (3.3 and 6.5 GHz) is greater (>20 ns). Owing to the relatively large number of obstacles, it indicates that the effective long multipath in the UMi scenario is more abundant. This result is consistent with the propagation characteristics of low-frequency signals, which exhibit a stronger diffraction ability. The UMi scene contains many obstacles, such as trees and buildings. Furthermore, the delay spread difference between the two frequency bands is minimal (<5 ns), indicating similar multipath delay characteristics. Therefore, the 5G system framework structure designed on the basis of channel characteristics of the 3.5 GHz band is applicable to the 6 GHz band in the 5G-Advanced system.
V-B THz Channel Measurements
@@
It is important to conduct extensive THz measurements to build a channel model for THz communication [146]. To date, many research organizations and institutions have conducted multiple THz channel measurement campaigns in the abovementioned frequency bands. We have classified and summarized some representative THz channel measurement campaigns by frequency bands, striving to comprehensively cover all currently measured THz frequency bands (100-1100 GHz), as shown in Table III.
@@
In the THz spectrum allocated for land mobile applications, as illustrated in Fig. 14, THz measurement campaigns are relatively concentrated around the frequencies of 100, 140, 220, and 300 GHz [147, 17, 31, 148, 149, 150, 151, 152, 153, 154, 155, 156, 157, 158, 159, 160, 161, 162, 163, 164, 165, 166, 167, 168, 169, 170, 171, 172, 173, 174, 175, 176, 177, 178, 179, 180, 181, 182, 183, 184, 185, 186, 187, 188, 189, 190, 191, 192, 193, 194, 167, 195, 196, 197, 198, 199, 200, 201, 202, 203, 178, 204, 205, 206, 207, 208, 209, 210, 211, 212]. The frequency of 100 GHz represents the upper limit of the ITU channel model. The primary objective of most studies is to investigate the applicability of existing channel models in the THz frequency band. Due to the non-linear factors of atmospheric absorption, the selection of actual frequencies primarily considers the “atmospheric window” frequency bands. In the THz frequency band, the “atmospheric window” refers to special frequency bands, such as 140 GHz and 220 GHz, where propagation attenuation is relatively low. Hence, these bands have been the focus of extensive research. Additionally, 300 GHz falls within the lower boundary of the THz frequency range defined by the ITU radio regulations, making it a frequency band of significant interest. However, in the 450-1100 GHz band [199, 213, 214, 215], there are fewer channel measurement campaigns limited by hardware capabilities, such as instruments and equipment. Due to limitations in transmitted power, the majority of studies focus only on reflection, scattering, and attenuation characteristics.
@@
Second, the measurement platforms in the summarized THz channel measurement campaigns are introduced. Most measurement platforms widely used in THz channel measurement campaigns are frequency domain vector network analyzer (VNA)-based platform [216]. In the VNA-based platform, the Tx and Rx are directly connected to the VNA. Owing to its simple structure, synchronized Tx and Rx, large bandwidth, and high dynamic range, this measurement platform is mainstream. However, due to the high loss of the cables, the measurement distance is usually short but can be extended to 100 m by radio-over-fiber (RoF) [178]. To conduct long-distance channel measurements, a correlation-based time-domain channel measurement platform is used, which operates with the strong autocorrelation of pseudo-random noise (PN) sequence. It has also been used in several THz channel measurements. Since the Tx and Rx are independent of each other and do not require cable connections, the platform is preferable for measurements in outdoor environments [148]. In addition, THz time-domain spectroscopy system (THz-TDS)-based [217] channel measurement platforms are limited by the output power and are commonly used for measuring the reflection and diffraction properties of materials.
@@
Third, the scenarios of the channel campaigns are presented. THz band channel measurements focus on desktops, meeting rooms, offices, data centers, computer motherboards, and indoor factory scenarios in indoor scenarios, and UMi, canyons, and streets in outdoor scenarios, covering vision scenarios, such as wireless local area networks (WLANs), wireless personal area networks (WPANs), data center networks (DCNs), and intradevice communication (IDCs) [218]. However, wireless backhaul/fronthaul scenarios have been poorly studied [219]. Most of the channel measurement campaigns in these scenarios are in the 100-450 GHz frequency band due to the limited hardware performance in the high-frequency bands. Furthermore, studies include channel characteristics, which will be described in detail in the next subsection.
TABLE III: Summary of main THz channel measurement campaigns
@@
132/1.2
Correlation-
based
Indoor factory, UMi
Cross-polarization discrimination,
PL, DS, K, AS
[211, 137]
@@
Indoor, outdoor rooftop
PL, SF, DS, snow attenuation
[149], [207]
140/1.536
Correlation-
based
Laboratory
PL, SF, K
@@
142/4
VNA
Entrance hall,
residential street
Power proportion
per reflection order
[212]
@@
corridor, lobby)
PL, SF, DS, AS, K,
cluster numbers,
intra/inter-cluster parameters
[186, 187],
[205]
@@
Table III (Continued): Summary of main THz channel measurement campaigns
Frequency
/BW (GHz)
Measurement
platforms
Scenarios Channel characteristics Reference
@@
V-C Important Channel Propagation Characteristics in THz Bands
Current research on THz characteristics can be categorized into two main types [31]. The first focuses on the material characteristics at THz bands, which actually have an impact on radio propagation (e.g., reflection) [220]. The second is dedicated to newly emerging channel characteristics in THz bands, such as atmospheric loss and channel sparsity.
V-C1 Reflection characteristics
Research on propagation mechanism characteristics is the key foundation and fundamental premise for THz channel characteristics research. Reflections are fundamental propagation mechanisms that should be thoroughly characterized to establish possible THz wireless links due to the short wavelength introduced in Section I. Therefore, this subsection focuses on the analysis and modeling of the reflection propagation characteristics of THz waves.
@@ V-C1 Reflection characteristics
At low frequencies, the surface of a material is considered primarily as a smooth surface. However, as the frequency increases to the THz band, the wavelength is short enough to be very close to the surface roughness of typical materials. Therefore, in the THz band, the material cannot be considered a smooth surface. The modified reflection coefficient is obtained by introducing a roughness factor $\rho$. The Rayleigh roughness factor can be expressed as [223]
@@
where $\sigma_{h}$ denotes the standard deviation height of the material surface roughness. The modified reflection coefficient $\Gamma^{\prime}$ is expressed as
@@
The modified reflection coefficient model is a function of the frequency and angle of incidence, and in reference [148], this phenomenon was verified by experiments. As an example, based on the modified Fresnel model, frequency-angle two-dimensional reflection coefficient modeling for vertical polarization is proposed. The model of non-metallic materials is written as:
@@
where $a$, $b$, $c$, $d$ are the parameters to be fitted, and their means are shown in [148]. The plasterboard results are shown in Fig. 17. The analysis revealed a low root-mean-square (RMS) error of approximately 0.1 for the glass, tile, board, and plasterboard materials, indicating a fit between the points of the measured data and the modified reflection coefficient model. The reflection coefficient of the material and its growth rate increased with increasing angle of incidence. Therefore, the model characterizes the reflection coefficient as a function of the frequency and angle of incidence.
@@
V-C2 Path loss
THz signals transmitted over channels have distinct characteristics compared to lower frequencies, leading to novel and urgent challenges [47]. Representatively, the spreading loss increases quadratically with frequency. Besides, with the THz band operating at exceptionally high frequencies, signals transmitted over a THz link are severely affected by the spreading loss [17]. The channel characteristics of high propagation loss can be characterized by path loss. A novel index is essential for characterizing and modeling path loss in the THz channel. The modeling of path loss in the THz band follows 5G modeling methods, such as the FI model, ABG model, and CI model shown in (1), (3), and (2), respectively, in Section II.A. Based on the references in Table III, the PLE of the CI model $\alpha_{\mathrm{CI}}$ for indoor and outdoor scenarios is summarized in Fig. 18. The PLE of the THz band is close to 2 in most cases, especially in LOS scenarios. In addition, the PLEs of the NLOS scenarios are markedly higher than those of the LOS scenario, even up to 4.6. This is because the blocking losses caused by walls, cars, etc., can be as large as tens of dB in NLOS scenarios, which means that severe attenuation can occur.
@@ V-C2 Path loss
Due to the wide span of the THz band, it will show a stronger frequency dependence. To further investigate the frequency dependence of path loss in THz bands, [17] used the ABG model to fit the measured path loss. The measured path loss and ABG model are plotted in Fig. 19. The frequency dependence factor $\gamma_{\mathrm{ABG}}$ and distance dependence factor $\alpha_{\mathrm{ABG}}$ are 2.1 and 1.93, respectively. This finding indicates that the frequency dependence is slightly stronger than that in the free space [47].
@@
Figure 19: Multi-frequency measurement data and path loss modeling results from 220 to 330 GHz [17].
V-C3 Atmospheric loss
In outdoor environments, THz signal propagation is strongly affected by atmospheric attenuation and weather-dependent scattering [224]. The primary component of atmospheric loss arises from molecular absorption by oxygen and water vapor, introducing both frequency-selective and distance-dependent loss characteristics into the channel [225]. Oxygen absorption is prominent in certain sub-100 GHz bands, such as the well-known 60 GHz peak, whereas water vapor absorption dominates above 100 GHz. For terrestrial deployments beyond 100 GHz, such as UMi and short-range backhaul links within 200 m, gas attenuation must be incorporated into the link budget. Specifically, for the 100-300 GHz band, the attenuation ranges from 0.106-1.293 dB at 200 m and 0.53-5.72 dB at 1,000 m. In the 300-450 GHz band, the corresponding values increase significantly, reaching 1.29-68.76 dB and 6.45-343.8 dB, respectively [226]. At higher altitudes, such as in airborne or satellite communications, attenuation becomes strongly altitude-dependent and typically decreases with elevation due to reduced pressure and temperature.
In addition to gaseous attenuation, various weather conditions, including rain, fog, and snow, introduce additional propagation variability. Rain causes relatively stable attenuation across frequencies and is well-characterized by existing droplet models, making it more predictable in system-level design [227, 228]. Fog exhibits a stronger frequency dependence and is more sensitive to ambient humidity and temperature. Snow presents the greatest uncertainty due to its highly variable particle shapes and the marked differences between wet and dry conditions [207]. These weather-induced effects can impact both signal strength and link stability, and thus must be carefully considered in the design of THz systems that target long-range applications such as vehicular, aerial, or backhaul communication links.
V-C4 Channel sparsity